# StableSketcher: Enhancing Diffusion Model for Pixel-based Sketch Generation via Visual Question Answering Feedback

### v1

@@ 5.4 User study
We conducted a ranking-based user study with 46 participants. For each prompt, participants anonymously compared five model outputs and ranked them from 1, indicating the best, to 5, indicating the worst, along three criteria: Sketch Characteristics, Prompt Fidelity, and Human-Drawn. Table X reports mean ranks, where a lower score indicates a stronger preference. Our method achieved the best mean rank on all criteria, with a score of 1.9 for Sketch Characteristics, 1.7 for Prompt Fidelity, and 1.7 for Human-Drawn, resulting in an overall score of 1.7. Compared to the strongest baseline in column d, which achieved 2.3 for Sketch Characteristics, 2.2 for Prompt Fidelity, 2.2 for Human-Drawn, and 2.2 overall, our method improved the mean rank by 17 percent for Sketch Characteristics and by 23 percent for both Prompt Fidelity and Human-Drawn, yielding a 23 percent relative gain in the overall mean rank. These results indicate consistent user preference for our sketches in terms of stylistic abstraction, textual prompt fidelity, and perceived human-likeness.
@@ 5.5 Ablation on VAE Loss Functions
We additionally considered binary cross-entropy (BCE) loss; however, its constraint that outputs lie in $[0,1]$ is incompatible with the LPIPS objective and our VAE decoder configuration. Optimizing with LPIPS alone yielded thick yet temporally consistent contours, whereas coupling L1 with LPIPS delivered smaller perceptual gains than the MSE–LPIPS pairing and left LPIPS effectively unchanged ($\approx 1\times 10^{-3}$) after 15 epochs. Taken together, the MSE–LPIPS combination offered the most favorable trade-off between pixel-level fidelity and perceptual sketch quality, as summarized in Figure 11.
@@ 6 Conclusion
We proposed StableSketcher, a framework that extends Stable Diffusion to human-drawn, pixel-based sketches by fine-tuning the latent autoencoder with a reconstruction–perceptual hybrid loss and introducing a reinforcement learning strategy guided by a VQA-based reward function. To support this, we introduced SketchDUO, the first dataset to provide triplets of sketch images, fine-grained captions, and QA pairs, enabling multimodal learning signals tailored to abstract sketch generation. In addition to this unique triplet structure, SketchDUO also incorporates positive examples that reflect desired abstraction and negative examples capturing common errors such as over-shading, excessive detail, or photorealistic bias, offering a contrastive design that enriches model training. Together, StableSketcher and SketchDUO enable fine-grained prompt alignment and better disentanglement of semantic correctness from stylistic faithfulness.
Experimental results confirmed the effectiveness of our approach. StableSketcher achieved the lowest FID (143.68) and the highest TIFAScore (0.68) across all configurations, outperforming Stable Diffusion v1.5 and v2.1 baselines. Compared to the baselines, StableSketcher reduced FID by up to 30.8% (v1.5) and 25.3% (v2.1), while improving TIFAScore by about 15 – 23% and CLIPScore by 4 – 10%, confirming consistent gains across metrics. In contrast, BERTScore showed little difference, suggesting that while it captures the overall semantic impression of an image, it fails to evaluate whether the generated output accurately and precisely reflects the input textual prompt—highlighting the limitation of caption-based metrics for measuring text-to-image generation prompt fidelity. Ablation studies further demonstrated that the proposed VQA-based reward provided more reliable improvements in prompt fidelity, especially for element-level conditions such as object count and background simplicity. Qualitatively, StableSketcher consistently generated sketches with clearer contour abstraction and reduced noise. A user study further validated its superiority, ranking StableSketcher highest in terms of sketch quality, semantic alignment, and perceived human-likeness.
Two limitations remain with respect to data coverage and evaluation. SketchDUO currently comprises 30 categories and 35.8K sketches, which constrains generalization to long-tail objects, multi-object scenes, and stylistic diversity. To address these limitations, we plan to expand the dataset in both breadth and depth by adding new classes, varying style factors (e.g., stroke density and thickness, contour versus hatching, perspective), and enriching annotations beyond captions and QA to include part labels, keypoints, and per-stroke metadata (order, length, curvature).

### v2

@@ 5.5 User study
Figure 9 shows the Google Forms-based interface used in our user study. The evaluation comprised three sections: Sketch Characteristics, Prompt Fidelity, and Human-Likeness. In the Sketch Characteristics section, participants assessed whether each image conformed to our definition of a sketch, with emphasis on simplicity, abstraction, black line drawing on a white background, and the absence of color, shading, and texture. In the Prompt Fidelity section, participants evaluated how well each image reflected the elements explicitly specified in the prompt, including object identity, attributes, and background conditions. In the Human-Likeness section, participants judged the extent to which each image resembled a human-drawn sketch rather than an overly synthetic or mechanically rendered output. Each section consisted of 10 questions. To ensure full coverage of SketchDUO, we randomly partitioned its 30 classes into three disjoint sets of 10 classes and assigned one set to each section. For each class, the image used in the user study was randomly sampled from the corresponding class instances. In this way, all 30 classes were included exactly once in the overall study, while the evaluated examples were randomly selected.
We conducted a ranking-based user study with 46 participants. For each question, participants compared five model outputs and assigned ranks from 1 to 5, where 1 indicated the most preferred result and 5 indicated the least preferred result under the corresponding criterion. Table 5 reports the mean rank for each method, where lower values indicate stronger user preference. Columns (a)–(d) correspond to Stable Diffusion v1.5, Stable Diffusion v2.1, Stable Diffusion v1.5 with U-Net fine-tuning, and Stable Diffusion v1.5 with both U-Net and VAE fine-tuning, respectively. Our method achieved the lowest mean rank across all criteria, obtaining 1.9 for Sketch Characteristics, 1.7 for Prompt Fidelity, and 1.7 for Human-Likeness, with an overall mean rank of 1.7. Among the baselines, the strongest performance was obtained by the model in column (d), which achieved mean ranks of 2.3, 2.2, and 2.2 on the three criteria, respectively, and 2.2 overall. Compared with this baseline, our method improved the mean rank by 0.4 for Sketch Characteristics and by 0.5 for both Prompt Fidelity and Human-Likeness, resulting in an overall improvement of 0.5. These results indicate that participants consistently preferred our outputs over the baselines in terms of sketch style, prompt alignment, and perceived human-likeness.
@@ 5.6 Ablation on VAE Loss Functions
We additionally considered binary cross-entropy (BCE) loss; however, its constraint that outputs lie in $[0,1]$ is incompatible with the LPIPS objective and our VAE decoder configuration. Optimizing with LPIPS alone yielded thick yet consistent contours, whereas coupling L1 with LPIPS delivered smaller perceptual gains than the MSE–LPIPS pairing and left LPIPS effectively unchanged ($\approx 1\times 10^{-3}$) after 15 epochs. Taken together, the MSE–LPIPS combination offered the most favorable trade-off between pixel-level fidelity and perceptual sketch quality, as summarized in Figure 13.
6 Discussion
The quantitative results demonstrate the effectiveness of our framework. StableSketcher achieved the lowest FID (143.68) and the highest TIFAScore (0.68) across all configurations, outperforming the Stable Diffusion v1.5 and v2.1 baselines. These gains indicate that the proposed reconstruction–perceptual hybrid loss and VQA-based reinforcement learning improve alignment between generated sketches and fine-grained textual prompts.
BERTScore showed negligible differences between our model and the baselines. This suggests that, although BERTScore captures the overall semantic impression of an image, it is less effective at evaluating whether specific element-level conditions in the prompt are faithfully reflected in the output. This observation further supports the need for our VQA-based reward function in improving prompt fidelity.
Another notable finding is that Stable Diffusion v1.5 outperformed v2.1 in sketch generation. While v2.1 is optimized for high-quality photorealistic synthesis through more aggressively filtered training data [2], this bias appears to hinder abstraction-oriented generation. By contrast, v1.5 was trained on a broader and less aggressively filtered distribution, including LAION-2B(en) and LAION-Aesthetics v2.5+ subsets [1], which may make it more adaptable to non-photorealistic domains. This distributional difference likely explains the stronger performance of the v1.5-based StableSketcher.
Despite these advances, our study still has several limitations. SketchDUO remains relatively small, containing only 30 categories and 4.7K sketches, and it also exhibits a Western-centric bias, which may limit generalization across broader domains and more diverse cultural contexts. In addition, our evaluation is conducted within the controlled style space defined by SketchDUO, and robustness to broader external human-drawn sketch distributions remains an important direction for future work. Our quantitative evaluation also relies on standard image-generation metrics, including FID, CLIPScore, and BERTScore. Although these metrics are widely used in recent diffusion-based sketch generation studies and support direct comparison with prior work, they are not specifically designed to capture sketch abstraction. To mitigate this limitation, we complemented the quantitative evaluation with a user study and are currently exploring dedicated evaluation metrics in follow-up work. Finally, the annotation pipeline would benefit from more scalable semi-automated strategies, which may improve annotation efficiency while enabling broader category coverage and more diverse sketch content. Although our current setup is not fully dependent on closed-source systems, GPT-4o was used only for initial caption drafting, whereas QA pairs were constructed using the open-source TIFA pipeline with LLaMA 2 and UnifiedQA. Reducing reliance on closed-source captioning models and developing stronger open-source alternatives remain important directions for future work.
@@ 7 Conclusion
In this work, we proposed StableSketcher, a framework for generating human-drawn, pixel-based sketches with Stable Diffusion, supported by SketchDUO, a novel dataset containing triplets of sketch images, fine-grained captions, and QA pairs. By incorporating positive examples of desired abstraction and negative examples of common errors, such as over-shading or photorealistic bias, our contrastive design helps disentangle semantic correctness from stylistic faithfulness.
Beyond its technical contributions, StableSketcher may support interactive sketch-generation applications. For example, it could support co-drawing with an artificial intelligence assistant for older adults or be used in creative drawing activities for children through iterative sketch expansion. More broadly, our work suggests the potential of sketch-oriented generative models for applications in education, therapy, and lifelong learning.
To address the limitations identified in this study, we are currently undertaking follow-up research. We are constructing a larger and more diverse dataset encompassing 300 categories to alleviate the scale and cultural constraints of the current dataset. At the same time, we are developing novel evaluation metrics designed to better capture sketch abstraction and stylistic fidelity beyond general-purpose image metrics. These efforts aim to provide a stronger foundation for sketch-oriented text-to-image generation.